# Spin Alignment of Vector Mesons Induced by Local Spin Density Fluctuations

### v2

@@ I Introduction
Understanding such measurements has been a challenge for theorists. Usually the fast rotation and the strong magnetic field created in non-central collisions have been considered as the main causes of spin polarization or alignment. It was found that the global rotation leads to global spin polarization of (anti-)hyperons while magnetic field can induce their splitting [7]. For spin alignment, it has been found that the global rotation leads to $\rho_{00}<1/3$ [8] while the magnetic field induces $\rho_{00}>1/3$ [9]. However, the global rotation or magnetic field in heavy ion collision is believed to be too weak to generate enough $|\rho_{00}-1/3|$ comparable to the experiment data [10, 11]. For example, consider the spin alignment contributed from the global spin polarization Λ/$\Lambda/\bar{\Lambda}$, which gives $\rho_{00}\approx\frac{1}{3}-\frac{4}{9}P_{\Lambda}P_{\bar{\Lambda}}$, and the global spin polarization contributes to the deviation of the magnitude $\frac{4}{9}P_{\Lambda}P_{\bar{\Lambda}}\approx 10^{-6}$, which is too small compared to the experimental data $\sim 10^{-2}$. Recently, Xin-li Sheng proposed that the vector meson field induced by the motion of $s/\bar{s}$ quark can contribute to spin alignment [12, 13, 14], and the authors of Ref.[15] found that the self-energy of quarks can modify the spin polarization and spin alignment. Besides, the global spin alignment has been understood in local axial charge currents fluctuation [16], the helicity polarization[17], the local vorticity and the anisotropic expansion of the fireball[11], the plasma fields [18], as well as light front quarks[19], and so on. The relation between the tensor polarization and quark spin correlations has been explored in [20]. However, the origin of spin alignment of vector mesons still remains unclear.
@@
On the other hand, spin alignment of vector mesons in most central collision 0-20% has been also observed in STAR [5], i.e., $\rho_{00}$ of $\phi$ is smaller than $1/3$ in high energy collision, while larger than $1/3$ at low energy. This observation is confusing from the traditional understanding based on the scenario of rotation and magnetic field, because at high energy and most central collisions, magnetic field, rotation and baryon chemical potential are supposed to be weak, thus $\rho_{00}=1/3$ is expected. This might indicate that the spin alignment of vector mesons is not coming from external rotation and magnetic field.
@@
Therefore, a question arises naturally: Is it possible that the spin alignments of vector mesons are generated by the intrinsic dynamics of QCD? In other words, we want to investigate if the spin alignment could be generated by the inner interaction of quarks, instead of the environment such as rotation or magnetic field. In this work, we propose a new mechanism for the spin alignment of vector meson: the spin density fluctuation of quarks. It is found that the spin of vector mesons depends on the local spin density of quarks but does not depend on the sign of the spin polarization, i.e., whether the number of spin-up quarks are more than that of spin-down quarks or vice versa, the same spin alignment can be obtained. Thus, we should pay more attention on the local spin polarization instead of the global spin polarization of quarks, because the local fluctuation of spin polarization vanishes after averaging over the system.
@@
Another key ingredient is the interactions between quarks. In principle, the spin structure of meson consist of contribution from both quark and gluons. However, it is still quite challenging to consider both quark and gluon in meson. Thus, we here consider model with quarks only, and the contribution of gluons can be regarded as the coupling constants of quark interactions. From the view point of four-fermion effective interactions, the scalar interaction is of greatest interest, for its close relation with the spontaneous chiral symmetry breaking in the vacuum and its restoration at finite temperature. However, it is worthy to point out that the axial-vector and tensor interactions are of importance for the study of spin alignment of vector mesons.
@@ II Global spin polarization under Mean Field Approximation
From the view of simple quark model, the spin of vector meson depends on the spins of its constituent quarks. And in the heavy ion collision, QGP is generated and as the fireball expands, its temperature decreases and mesons are emitted at freeze-out temperature from a local equilibrium region, the spin density, i.e., the density of spin-up minus the density of spin-down, of quark and anti-quark determines the distribution of emitted mesons with different spins. Thus, we firstly need to find appropriate quantities to describe the spin density. Below we will show that the axial-vector operator is related to the spin polarization(SP) while tensor operator to the magnetic moment polarization(MMP), both are important to describe the spin alignment of vector mesons.
@@ II.1 Axial-Vector and Tensor operator
Consider the Axial-Vector operator $\bar{\psi}\gamma^{\mu}\gamma^{5}\psi$ for free quark of one flavor, here we take the $z$-axis without loss of generality, i.e., $\mu=3$. With the standard procedure of quantum field theory, use the form of quark field expanded in momentum space, we could easily find that the integral of Axial-Vector operator over all space could be divided into four terms:
@@
where we define the abbreviation
@@
in the small momentum limit, i.e., $\mathbf{p}\rightarrow 0$, or large mass limit, $J^{\uparrow}_{+}$ is the total number of quark with spin-up, $N_{+}^{\uparrow}=\int_{p}a_{p}^{\uparrow\dagger}a_{p}^{\uparrow}$, and the meaning of $\delta J$ is clear, the net-spin angular momentum(up to a factor /2Planck-constant-over-$\hbar/2$). And at medium or large momentum, $J^{\uparrow}_{+}$ is not exact the number of net-spins, however, we can still relate it to the number of net-spins:
@@
and $c$ is a function of temperature and quark mass, and in a equilibrium system, for example, $c\approx 0.22$ for $T=0.15$GeV and $m_{q}=0.33$GeV. And $\delta J$ can be regarded as spin polarization(SP).
Similar to the axial-vector interaction, tensor interaction of quarks will also cause spin polarization of $\phi$ meson. The tensor operator $\bar{\psi}\Sigma^{3}\psi=\bar{\psi}i\gamma^{1}\gamma^{2}\psi$ indicates the imbalance between magnetic-moment-up and -down, i.e., magnetic moment polarization(MMP), instead of spin polarization in the case of axial-vector operator, which can be seen from:
@@ II.2 Global Quark Spin Polarization
The simplest case for spin polarization is a global and homogeneous spin polarization, and it is straightforward to consider that the SP is induced spontaneously from the quark interactions, where a global and homogeneous background, i.e., condensate, can be obtained. Here we assume a NJL-type four-fermion interaction of axial-vector channel, $G_{A}(\bar{\psi}\gamma^{\mu}\gamma^{5}\psi)^{2}$, where $\mu=3$ is taken in the following. All the following argument can be applied to the case of tensor interaction, $G_{T}(\bar{\psi}\Sigma^{3}\psi)^{2}$, and we neglect this case here. Then the Lagrangian for quarks with NJL-type scalar and axial-vector interactions has form of:
@@
where $m_{0}$ is current quark mass. Similar to the definition of chiral condensate $\langle\bar{\psi}\psi\rangle$, we here also define the axial-vector condensate $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$, then we can write down the new Lagrangian under mean field approximation(MFA) as usually used in NJL model [21, 22]:
@@
then the thermodynamic potential
@@
Usually, the chiral condensate is not zero at low temperature, $\langle\bar{\psi}\psi\rangle\neq 0$, similarly, as discussed above, a nonzero tensor condensate, i.e., $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle\neq 0$, is expected. However, it is found numerically that for $G_{A}\lesssim 2G_{S}$, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ is always zero at finite temperature. And for $G_{A}\gtrsim 2G_{S}$, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ could be nonzero, however, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ and $\langle\bar{\psi}\psi\rangle$ cannot be nonzero simultaneously, in this case where $m_{q}=0$ and $a\neq 0$, only spin-up of quarks exists, which leads to $\rho_{00}\sim 1$ of vector mesons.
@@
In a equilibrium and none-polarized system, $\rho_{00}=1/3$ is obvious. However, due to the global spin polarization or spin fluctuation or any other possible reasons, $N^{\uparrow}_{s/\bar{s}}\neq N^{\downarrow}_{s/\bar{s}}$ thus $\rho_{00}\neq 1/3$ could happen. Define $N_{s}=N^{\uparrow}_{s}+N^{\downarrow}_{s}$, which is the total density of $s$ quark with both spin-up and spin-down, and $\delta N_{s}=N^{\uparrow}_{s}-N^{\downarrow}_{s}$, which is the number of net-spin of $s$ quark. Assume that $\delta N_{s}$ is small, i.e., $\delta N_{s}/N_{s}\ll 1$, then we obtain:
@@
it is clear that the spin polarization of $s/\bar{s}$ quark directly lead to spin alignment of $\phi$ meson, and the deviation of $\rho_{00}$ from $1/3$, i.e., $\delta\rho_{00}=\rho_{00}-1/3$, could either be positive or negative, which depends on the correlation of net-spin of $s$ quark and $\bar{s}$ quark. Similar form also can be found in [20]. In a local small region, for example, if spin of $s$ quark tends to be same with $\bar{s}$, let’s say induced by global rotation, then $\langle\delta N_{s}\delta N_{\bar{s}}\rangle>0$, thus $\delta\rho_{00}<1/3$; If magnetic field is presented, then $\langle\delta N_{s}\delta N_{\bar{s}}\rangle<0$ is obtained for $s$ and $\bar{s}$ has opposite charge, thus $\delta\rho_{00}>1/3$.
It is worthy to point out that Here we don’t consider the difference between $s$ and $\bar{s}$, i.e., at $\mu_{B}=0$. And none-zero baryon chemical potential lead larger deviation of $\rho_{00}$ from $1/3$, however, here we will not make further discussion. Then consider another vector meson $K^{*0}$, which is composed of $d$ and $\bar{s}$, and its spin alignment has a similar form:
@@
Figure 2: Schematic diagram of quark spin fluctuation. (a) No global and local spin polarization. However, the local spin polarization of quark could be none-zero for a quantum system at finite temperature due to the fluctuation even no global spin polarization, as shown in (b).
Although none-zero $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ can not be obtained spontaneously with the axial-vector interaction, as we have proved in previous section, $\bar{\psi}\gamma^{3}\gamma^{5}\psi$ is still not exactly zero at any local region of the system due to the fluctuation nature of a quantum system at finite temperature, as illustrated in FIG.2. On the other hand, the the spin alignment $\delta\rho_{00}$ doesn’t require a none-zero global and homogeneous $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ but the spin correlation $\langle\delta N_{s}\delta N_{\bar{s}}\rangle\neq 0$. Consider the free quark gas where the interaction between quarks is ignored, then the fluctuation of quark and anti-quark is completely independent, i.e., $\langle\delta N_{s}\delta N_{\bar{s}}\rangle=\langle\delta N_{s}\rangle\langle\delta N_{\bar{s}}\rangle=0$, thus no spin alignment of vector meson will appear. If we turn on the quark interaction, let’s say a attraction between spins, then the spin of quark as well as anti-quark tend to point to the same direction in a local region, i.e., $\langle\delta N_{s}\delta N_{\bar{s}}\rangle>0$, which means the sign of local spin polarization, $\delta N$, doesn’t affect the sign of spin alignment , $\delta\rho_{00}$, and the contribution from local spin polarization will not be washed by the average over all events or all the system.
@@ III.2 Spin Alignment of $\phi$
which relates to the axial-vector and the tensor operators, and that’s why we consider these two operators. Now, we show how to calculate $\langle\delta J^{2}\rangle$, and the process is total same for $\langle\delta M^{2}\rangle$. Consider quarks in a spin polarization background where a none-zero spin potential $a$ is presented, then define the thermal dynamic potential:
@@
[…]
@@
and now the partition function is a function of $a$. Then we consider the second derivative of $\Omega$ with respect to $a$, which is found to have two terms:
@@
which is zero for $a=0$, and as we will also prove that $\langle\delta J\rangle=0$ even with axial-vector interaction but small coupling constant $G_{A}$, and the second term:
@@
If we take $a\rightarrow 0$, then the first term will be zero because $\lim_{a\rightarrow 0}\langle N_{s}\rangle=0$, while the second term gives none-zero contribution, which is exactly what we want:
$\langle(\delta J)^{2}\rangle=-VT\frac{\partial^{2}}{\partial a^{2}}\Omega(a)|_{a\rightarrow 0},$ (26)
@@
Then consider the thermodynamics potential, the Feynman diagram of which is shown in FIG.3 and only the leading order of axial-vector interaction with coupling constant $G_{A}$ is considered. The first one is of order $N_{c}^{2}G_{A}$ while the second one $N_{c}G_{A}$, where $N_{c}$ is the number of color. And similar with the case of chiral condensate, we only consider the first Feynman diagram, which is of form:
Figure 3: The thermal dynamics potential in the leading order with respect to the axial-vector interaction. Diagram (a) is of order $G_{A}N_{c}^{2}$ while (b) is $G_{A}N_{c}$.
@@
where $f(E)=1/(\text{e}^{E/T}+1)$ is Fermi-Dirac distribution. Now there is no $a$-independent term, and this is reasonable because that there is of course no spin polarization at zero spin potential $a=0$. Now we obtain the thermal dynamical potential:
@@
$\langle\delta J^{2}\rangle=32N_{c}^{2}G_{A}VTL^{2},$ (33)
where $V$ and $T$ are the volume and temperature of the fireball created in the heavy ion collision. The process for calculating $\langle(\delta M)^{2}\rangle$ is same, however, notice:

### v3

@@ I Introduction
Understanding such measurements has been a challenge for theorists. Usually the fast rotation and the strong magnetic field created in non-central collisions have been considered as the main causes of spin polarization or alignment. It was found that the global rotation leads to global spin polarization of (anti-)hyperons while magnetic field can induce their splitting [7]. For spin alignment, it has been found that the global rotation leads to $\rho_{00}<1/3$ [8] while the magnetic field induces $\rho_{00}>1/3$ [9]. However, the global rotation or magnetic field in heavy ion collision is believed to be too weak to generate enough $|\rho_{00}-1/3|$ comparable to the experiment data [10, 11]. For example, consider the spin alignment contributed from the global spin polarization Λ/$\Lambda/\bar{\Lambda}$, which gives $\rho_{00}\approx\frac{1}{3}-\frac{4}{9}P_{\Lambda}P_{\bar{\Lambda}}$, however, it was found $\frac{4}{9}P_{\Lambda}P_{\bar{\Lambda}}\approx 10^{-6}$, which is too small compared to the experimental data $\sim 10^{-2}$. Recently, Xin-li Sheng proposed that the vector meson field induced by the motion of $s/\bar{s}$ quark can contribute to spin alignment [12, 13, 14], and the authors of Ref.[15] found that the self-energy of quarks can modify the spin polarization and spin alignment. Besides, the global spin alignment has been understood in local axial charge currents fluctuation [16], the helicity polarization[17], the local vorticity and the anisotropic expansion of the fireball[11], the plasma fields [18], as well as light front quarks[19], and so on. The relation between the tensor polarization and quark spin correlations has been explored in [20]. However, the origin of spin alignment of vector mesons still remains unclear.
@@
On the other hand, spin alignment of vector mesons has also been observed in the most central collisions 0-20% at STAR [5], i.e., $\rho_{00}$ of $\phi$ is smaller than $1/3$ in high energy collision, while larger than $1/3$ at low energy. This observation is confusing from the traditional understanding based on the scenario of rotation and magnetic field, because at high energy and most central collisions, magnetic field, rotation and baryon chemical potential are supposed to be weak, thus $\rho_{00}=1/3$ is expected. This might indicate that the spin alignment of vector mesons is not coming from external rotation and magnetic field.
@@
Therefore, a question arises naturally: Is it possible that the spin alignments of vector mesons are generated by the intrinsic dynamics of QCD? In other words, we want to investigate if the spin alignment could be generated by the inner interaction of quarks, instead of the environment such as rotation or magnetic field. In this work, we propose a new mechanism for the spin alignment of vector meson: the spin density fluctuation of quarks. It is found that the spin of vector mesons depends on the local spin density of quarks but not on the sign of spin polarization, which means that the same spin alignment can be achieved regardless of whether there are more spin-up quarks than spin-down quarks, or vice versa.Thus, we should pay more attention on the local spin polarization instead of the global spin polarization of quarks, because the local fluctuation of spin polarization vanishes after averaging over the system.
@@
Another key ingredient is the interactions between quarks. In principle, the spin structure of a meson consists of contributions from both quarks and gluons. However, it is still quite challenging to consider both quarks and gluons in a meson simultaneously. Thus, we here consider model with quarks only, and the contribution of gluons can be regarded as the coupling constants of quark interactions. From the view point of four-fermion effective interactions, the scalar interaction is of greatest interest, for its close relation with the spontaneous chiral symmetry breaking in the vacuum and its restoration at finite temperature. However, it is worthy to point out that the axial-vector and tensor interactions are of importance for the study of spin alignment of vector mesons.
@@ II Global spin polarization under Mean Field Approximation
From the view of simple quark model, the spin of vector meson depends on the spins of its constituent quarks. And in the heavy ion collision, QGP is generated and as the fireball expands, its temperature decreases and mesons are emitted at freeze-out temperature from a local equilibrium region, the spin density, i.e., the density of spin-up minus the density of spin-down, of quark and anti-quark determines the distribution of emitted mesons with different spins. Thus, we firstly need to find appropriate quantities to describe the spin density. In the following, we will show that the axial-vector operator is related to the spin polarization(SP) while tensor operator to the magnetic moment polarization(MMP), both are crucial for describing the spin alignment of vector mesons.
@@ II.1 Axial-Vector and Tensor operator
Consider the axial-vector operator $\bar{\psi}\gamma^{\mu}\gamma^{5}\psi$ for free quark of one flavor, here we take the $z$-axis without loss of generality, i.e., $\mu=3$. With the standard procedure of quantum field theory and use the form of quark field expanded in momentum space, we could easily find that the integral of axial-vector operator over all space could be divided into four terms:
@@
where $N_{c}=3$ is number of color and we define the abbreviation:
@@
and in the small momentum limit, i.e., $\mathbf{p}\rightarrow 0$, or large mass limit, $J^{\uparrow}_{+}=N_{c}\int_{p}a_{p}^{\uparrow\dagger}a_{p}^{\uparrow}$ is the total number of quark with spin-up, and the meaning of $\delta J$ is clear: the net-spin angular momentum(up to a factor /2Planck-constant-over-$\hbar/2$). And at medium or large momentum, $J^{\uparrow}_{+}$ is not exact the number of net-spins, however, we can still relate it to the number of net-spins:
@@
and $c$ is a function of temperature and quark mass, and in a equilibrium system, for example, $c\approx 0.22$ for $T=0.15$GeV and $m_{q}=0.33$GeV. Thus, $\delta J$ can be regarded as spin polarization.
Similarly, the tensor operator $\bar{\psi}\Sigma^{3}\psi=\bar{\psi}i\gamma^{1}\gamma^{2}\psi$ indicates the imbalance between magnetic-moment-up and -down, i.e., magnetic moment polarization(MMP), instead of spin polarization in the case of axial-vector operator, which can be seen from:
@@ II.2 Global Quark Spin Polarization
The simplest case for spin polarization is a global and homogeneous spin polarization, and it is straightforward to consider that the SP is induced spontaneously from the quark interactions, where a global and homogeneous background, i.e., condensate, can be obtained. Here we assume a NJL-type four-fermion interaction of axial-vector channel, $G_{A}(\bar{\psi}\gamma^{\mu}\gamma^{5}\psi)^{2}$, where $\mu=3$ is taken in the following. All the following argument can be applied to the case of tensor interaction, $G_{T}(\bar{\psi}\Sigma^{3}\psi)^{2}$, and we neglect this case here. Then the Lagrangian for quarks with NJL-type scalar and axial-vector interactions has form of [21, 22]:
@@
where $m_{0}$ is current quark mass. Similar to the definition of chiral condensate $\langle\bar{\psi}\psi\rangle$, we here also define the axial-vector condensate $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$, then we can write down the new Lagrangian under mean field approximation(MFA) as usually used in NJL model:
@@
consider a quark system at finite temperature $T$ and volume $V$, and the grand canonical ensemble is usually used to describe the fireball in heavy ion collision[23], and more details on finite temperature field theory can be found in the textbook[24]. Then the thermodynamic potential
@@
Usually, the chiral condensate is not zero at low temperature, $\langle\bar{\psi}\psi\rangle\neq 0$, similarly, as discussed above, a non-zero tensor condensate, i.e., $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle\neq 0$, is expected. However, it is found numerically that for $G_{A}\lesssim 2G_{S}$, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ is always zero at finite temperature. And for $G_{A}\gtrsim 2G_{S}$, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ could be non-zero, however, $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ and $\langle\bar{\psi}\psi\rangle$ cannot be non-zero simultaneously, in this case where $m_{q}=0$ and $a\neq 0$, only spin-up of quarks exists, which leads to $\rho_{00}\sim 1$ of vector mesons.
@@
In a equilibrium and non-polarized system, $\rho_{00}=1/3$ is obvious. However, due to the global spin polarization or spin fluctuation or any other possible reasons, $N^{\uparrow}_{s/\bar{s}}\neq N^{\downarrow}_{s/\bar{s}}$ thus $\rho_{00}\neq 1/3$ could happen. Define $N_{s}=N^{\uparrow}_{s}+N^{\downarrow}_{s}$, which is the total density of $s$ quark with both spin-up and spin-down, and $\delta N_{s}=N^{\uparrow}_{s}-N^{\downarrow}_{s}$, which is the number of net-spin of $s$ quark. Assume that $\delta N_{s}$ is small, i.e., $\delta N_{s}/N_{s}\ll 1$, then we obtain:
@@
it is clear that the spin polarization of $s/\bar{s}$ quark directly lead to spin alignment of $\phi$ meson, and the deviation of $\rho_{00}$ from $1/3$, i.e., $\delta\rho_{00}=\rho_{00}-1/3$, could either be positive or negative, which depends on the correlation of net-spin of $s$ quark and $\bar{s}$ quark. Similar formula also can be found in [20]. In a local small region, for example, if spin of $s$ quark tends to be same with $\bar{s}$, let’s say induced by global rotation, then $\langle\delta N_{s}\delta N_{\bar{s}}\rangle>0$, thus $\delta\rho_{00}<1/3$; If magnetic field is presented, then $\langle\delta N_{s}\delta N_{\bar{s}}\rangle<0$ is obtained because that $s$ and $\bar{s}$ has opposite charge, thus $\delta\rho_{00}>1/3$.
It is worthy to point out that here we don’t consider the difference between $s$ and $\bar{s}$, i.e., at $\mu_{B}=0$. And we have found that non-zero baryon chemical potential leads to larger deviation of $\rho_{00}$ from $1/3$, however, here we will not make further discussion. Then consider another vector meson $K^{*0}$, which is composed of $d$ and $\bar{s}$, and its spin alignment has a similar form:
@@
Figure 2: Schematic diagram of quark spin fluctuation. (a) No global and local spin polarization. However, the local spin polarization of quark could be non-zero for a quantum system at finite temperature due to the fluctuation even no global spin polarization, as shown in (b).
Although non-zero $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ can not be obtained spontaneously with the axial-vector interaction, as we have proved in previous section, $\bar{\psi}\gamma^{3}\gamma^{5}\psi$ is still not exactly zero at any local region of the system due to the fluctuation nature of a quantum system at finite temperature, as illustrated in FIG.2. On the other hand, the spin alignment $\delta\rho_{00}$ doesn’t require a non-zero global and homogeneous $\langle\bar{\psi}\gamma^{3}\gamma^{5}\psi\rangle$ but the spin correlation $\langle\delta N_{s}\delta N_{\bar{s}}\rangle\neq 0$. Consider the free quark gas where the interaction between quarks is ignored, then the fluctuation of quark and anti-quark is completely independent, i.e., $\langle\delta N_{s}\delta N_{\bar{s}}\rangle=\langle\delta N_{s}\rangle\langle\delta N_{\bar{s}}\rangle=0$, thus no spin alignment of vector meson will appear. If we turn on the quark interaction, let’s say a attraction between spins, then the spin of quark as well as anti-quark tend to point to the same direction in a local region, i.e., $\langle\delta N_{s}\delta N_{\bar{s}}\rangle>0$, which means the sign of local spin polarization, $\delta N$, doesn’t affect the sign of spin alignment, $\delta\rho_{00}$, and the contribution from local spin polarization will not be washed out by averaging over all events or the entire system.
@@ III.2 Spin Alignment of $\phi$
which relates to the axial-vector and the tensor operators, and that’s why we consider these two operators. Now, we show how to calculate $\langle\delta J^{2}\rangle$, and the process is total same for $\langle\delta M^{2}\rangle$. Consider quarks in a spin polarization background where a non-zero spin potential $a$ is presented, then define the thermodynamic potential:
@@
where $Z(a)$ is the partition function at finite temperature with effective quark mass $m$ and the effective spin potential $a$ as well as the axial-vector interaction $G_{A}(\bar{\psi}\gamma^{3}\gamma^{5}\psi)^{2}$, which has form of
@@
where $\tau=it$ is imaginary time, and the integral over $\tau$ is from 0 to $\beta$, where $\beta=1/T$ [24]. And now the partition function is a function of $a$. Then we consider the second derivative of $\Omega$ with respect to $a$, which is found to have two terms:
@@
which is zero for $a=0$, and the second term:
@@
If we take $a\rightarrow 0$, then the first term will be zero because $\lim_{a\rightarrow 0}\langle N_{s}\rangle=0$, while the second term gives non-zero contribution, which is exactly what we want:
$\langle(\delta J)^{2}\rangle=-VT\frac{\partial^{2}}{\partial a^{2}}\Omega(a)|_{a\rightarrow 0}.$ (26)
@@
Then consider the thermodynamic potential, the Feynman diagram of which is shown in FIG.3 and only the leading order of axial-vector interaction with coupling constant $G_{A}$ is considered. The first one is of order $N_{c}^{2}G_{A}$ while the second one $N_{c}G_{A}$. And similar with the case of chiral condensate, we only consider the first Feynman diagram, which is of form:
Figure 3: The thermodynamic potential in the leading order with respect to the axial-vector interaction. Diagram (a) is of order $G_{A}N_{c}^{2}$ while (b) is $G_{A}N_{c}$.
@@
where $f(E)=1/(\text{e}^{E/T}+1)$ is Fermi-Dirac distribution, $p_{\perp}$ is the momentum in the $x-y$ plane while $E=\sqrt{p_{\perp}^{2}+p_{3}^{2}+m^{2}}$ is energy. Now there is no $a$-independent term, and this is reasonable because that there is of course no spin polarization at zero spin potential $a=0$. Now we obtain the thermodynamic potential:
@@
$\langle\delta J^{2}\rangle=32N_{c}^{2}G_{A}VTL^{2}.$ (33)
The process for calculating $\langle(\delta M)^{2}\rangle$ is same, however, notice: